# Exact Solution of Bipartite Fluctuations in One-Dimensional Fermions

### v1

@@ Abstract
Emergence of hydrodynamics in quantum many-body systems has recently garnered growing interest. The recent experiment of ultracold atoms [J. F. Wienand et al., arXiv:2306.11457] studied emergent hydrodynamics in hard-core bosons using a bipartite fluctuation, which quantifies how the particle number fluctuates in a subsystem. In this Letter, we theoretically study the variance of a bipartite fluctuation in one-dimensional noninteracting fermionic dynamics starting from an alternating state, deriving the exact solution of the variance and its asymptotic linear growth law for the long-time dynamics. To compare the theoretical prediction with the experiment, we generalize our exact solution by incorporating the incompleteness of the initial alternating state, deriving the general linear growth law analytically. We find that it quantitatively describes the experimentally observed variance growth without any fitting parameters.
@@
Introduction.— Relaxation to an equilibrium state is ubiquitous in quantum many-body dynamics, bringing up fundamental and intriguing questions, e.g., how an isolated quantum system relaxes to a thermal equilibrium state Polkovnikov et al. (2011); Eisert et al. (2015); Nandkishore and Huse (2015); Torres-Herrera et al. (2015); Luca D’Alessio and Rigol (2016); Mori et al. (2018); Abanin et al. (2019); Gong and Hamazaki (2022). Over decades, such quantum thermalization has been intensively studied from broad perspectives, such as the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994), integrability Kinoshita et al. (2006); Gring et al. (2012); Cazalilla (2006); Rigol et al. (2006, 2007); Calabrese et al. (2011); Fagotti and Essler (2013); Langen et al. (2015); Ilievski et al. (2015); Tang et al. (2018), many-body localization Basko et al. (2006); Žnidarič et al. (2008); Pal and Huse (2010); Bardarson et al. (2012); Serbyn et al. (2013); Huse et al. (2014); Schreiber et al. (2015); Lüschen et al. (2017); Kohlert et al. (2019), and quantum scars Bernien et al. (2017); Shiraishi and Mori (2017); Turner et al. (2018a, b); Lin and Motrunich (2019); Shibata et al. (2020). To date, many experiments involving ultracold atoms and trapped ions have observed various phenomena related to quantum thermalization Kinoshita et al. (2006); Gring et al. (2012); Trotzky et al. (2012); Langen et al. (2013); Schreiber et al. (2015); Langen et al. (2015); Clos et al. (2016); Kaufman et al. (2016); Neill et al. (2016); Lüschen et al. (2017); Bernien et al. (2017); Tang et al. (2018); Kohlert et al. (2019). For example, S. Trotzky et al. realized an isolated bosonic system on a one-dimensional lattice, observing the relaxation dynamics starting from an alternating state where the bosons occupy every other site Trotzky et al. (2012).
Recently, beyond the conventional quantum thermalization, hydrodynamic description based on local equilibrium has rapidly developed in quantum many-body systems. This situation is exemplified by the recent observation of electron fluids in strongly interacting systems Moll et al. (2016); Bandurin et al. (2016); Gooth et al. (2018); Krishna Kumar et al. (2017); Sulpizio et al. (2019); Aharon-Steinberg et al. (2022) and the success of generalized hydrodynamics being valid for integrable quantum models Castro-Alvaredo et al. (2016); Bertini et al. (2016); Bulchandani et al. (2017, 2018); Doyon and Yoshimura (2017); Doyon et al. (2017, 2018); Collura et al. (2018); De Nardis et al. (2018); Schemmer et al. (2019); Doyon (2020); Alba et al. (2021); Malvania et al. (2021); Bouchoule and Dubail (2022); Essler (2023). When considering such hydrodynamic description, it is of essence to scrutinize how a system approaches a local equilibrium state. One of the most useful quantities for diagnosing local equilibrium is a bipartite fluctuation that quantifies how a physical quantity in a subsystem fluctuates. In a recent experiment of ultracold atoms for hard-core bosons Wienand et al. (2023), J. F. Wienand et al. utilized the bipartite fluctuation of the particle number in quench dynamics starting from the alternating state, studying the local equilibrium and the emergence of fluctuating hydrodynamics in the quantum many-body system.
Despite the recent strong interest in emergent hydrodynamics in a quantum regime, exact and analytical results of the bipartite fluctuation with the initial alternating state have been elusive even for the noninteracting fermions Groha et al. (2018); com , although there are several numerical works Fujimoto et al. (2020); Jin et al. (2020); Fujimoto et al. (2021, 2022); Cecile et al. (2024); Aditya and Roy (2024); Bhakuni and Lev (2023). Hence, its fundamental and analytical understanding is highly desirable now, and it is of great importance to explain the recently observed bipartite fluctuation Wienand et al. (2023) quantitatively via analytical solutions.
@@
Figure 1: Schematic illustration of the main result. (a) One-dimensional fermions addressed in this work. The blue circles represent fermions occupying every other site at the initial time. This initial state is referred to as an alternating state. Under unitary time evolution, the fermions hop to neighboring sites. The quantity of our interest is the variance of the particle number in a subsystem $A$ for a time-evolved state. (b) Schematic for the variance growth. The abscissa and ordinate, respectively, denote time $t$ and a subtracted variance, which is defined by the variance from which one subtracts its initial value. The two solid lines represent our analytical results, namely $2t/\pi$ and $2(n_{\rm even}-n_{\rm odd})^{2}t/\pi$, obtained by considering the variances growing from the complete and incomplete alternating states. Here, $n_{\rm even}$ and $n_{\rm odd}$ are, respectively, averaged initial particle numbers at even and odd sites, quantifying the incompleteness of the initial alternating state.
@@
In this Letter, we consider one-dimensional fermions starting from the alternating state, theoretically studying time evolution for the variance of the particle numbers in a subsystem as schematically shown in Fig. 1 (a) and compare our theoretical prediction with the recent experimental data of Ref. Wienand et al. (2023). First, in the noninteracting fermions, we exactly compute infinite series involving the $n$th-order Bessel function $J_{n}(x)$ of the first kind, obtaining the exact and simple expression of the variance. We apply asymptotic analysis to the exact result, analytically deriving the linear growth law of the variance and its sub-leading correction for the long-time dynamics. Second, we compare our analytical result with the recent experiment of Ref. Wienand et al. (2023). For this purpose, we focus on the incompleteness of the initial alternating state realized in the experiment. We incorporate the incompleteness into our exact solution, deriving the general linear growth law of the variance. This general law can quantitatively explain the variance growth observed in Ref. Wienand et al. (2023) without any fitting parameters. The impact of the incompleteness is schematically displayed in Fig. 1 (b). Finally, we numerically investigate interaction effects on the variance growth using the time-evolving decimation (TEBD) method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019), finding that the presence of the interaction substantially alters the behavior of the variance.
@@
Before ending the introduction, it is worth noting that a bipartite fluctuation for a conserved quantity is related to an integrated current, which has been studied in nonequilibrium statistical mechanics Antal et al. (2008); Eisler and Rácz (2013); Moriya et al. (2019); Gamayun et al. (2020); Jin et al. (2021); Gopalakrishnan and Vasseur (2023); Wei et al. (2022). Previous literature has focused mainly on dynamics starting from a domain-wall initial state, studying a bipartite fluctuation by connecting it to an integrated current via a continuity equation. One of the notable results is the logarithmic growth of the variance in noninteracting fermions Antal et al. (2008).
@@
with the particle-number operator $\hat{n}_{j}\coloneqq\hat{a}_{j}^{\dagger}\hat{a}_{j}$ and the interaction parameter $U$. Here, the boundary condition is basically a periodic boundary condition ($\hat{a}_{L}=\hat{a}_{-L}$), but we use an open boundary condition ($\hat{a}_{L}=0$) in several cases. In the latter case, we will explicitly specify the boundary condition. The initial state is the alternating state defined by |$\ket{\psi_{\rm alt}}\coloneqq\prod_{j=-L/2}^{L/2-1}\hat{a}_{2j+1}^{\dagger}\ket{0}$. We denote a quantum state at time $t$ by |$\ket{\psi(t)}$ and assume that the state obeys the Schrödinger equation ${\rm i}d\ket{\psi(t)}/dt=\hat{H}\ket{\psi(t)}$. Here, the Dirac constant $\hbar$ is set to be unity.
In this Letter, we shall study the fluctuation of the particle number in a subsystem. The operator for the bipartite particle-number is defined by $\hat{N}_{M}\coloneqq\sum_{j=0}^{M-1}\hat{a}_{j}^{\dagger}\hat{a}_{j}$ with a positive integer $M$. Then, a generating function for the $n$th moment of $\hat{N}_{M}$ is given by $G_{M}(\lambda,t)\coloneqq\braket{e^{\lambda\hat{N}_{M}}}_{t}$, where we introduce the notation $\braket{\bullet}_{t}\coloneqq{\rm Tr}[\hat{\rho}(t)\bullet]$ with the density matrix $\hat{\rho}(t)\coloneqq\ket{\psi(t)}\bra{\psi(t)}$ |. We can obtain the moments by differentiating $G_{M}(\lambda,t)$ with respect to $\lambda$.
Exact solution for the variance of the noninteracting fermions.— We study the time evolution for the variance of $\hat{N}_{M}$ in the noninteracting fermions ($U=0$) under the thermodynamic limit ($L\rightarrow\infty$). As described in Sec. I of the Supplemental Material SM , the generating function $G_{M}(\lambda,t)$ becomes
@@
The detailed derivation of Eq. (4) is given in Sec. II of the Supplemental Material SM . The next task is to take the limit $M\rightarrow\infty$ in Eq. (4). As proved in Sec. III of the Supplemental Material SM , we have $\lim_{M\rightarrow\infty}M\left(1-J_{0}(4t)^{2}-2\sum_{k=1}^{M-1}J_{k}(4t)^{2}\right)/4=0$ for $t>0$ and $\lim_{M\rightarrow\infty}\sum_{k=1}^{M-1}kJ_{k}(4t)^{2}/2=4t^{2}\left(J_{0}(4t)^{2}+J_{1}(4t)^{2}\right)-tJ_{0}(4t)J_{1}(4t)$. We utilize these formulae and Eq. (4), deriving the following exact and simple expression of the variance $\sigma(t)^{2}\coloneqq\lim_{M\rightarrow\infty}\sigma_{M}(t)^{2}$,
@@
The detailed derivation of Eq. (6) is given in Sec. IV of the Supplemental Material SM . The essential consequence of Eq. (6) is that the variance is proportional to time $t$ for $t\gg 1$. This linear growth law is entirely different from the variance growth in noninteracting fermionic dynamics starting from the domain-wall state |$\ket{\psi_{\rm DW}}\coloneqq\prod_{j\in\mathbb{Z}_{<0}}\hat{a}_{j}^{\dagger}\ket{0}$, where the variance was shown to obey the logarithmic growth Antal et al. (2008); Moriya et al. (2019). Note that the linear growth law was numerically confirmed in Ref. Fujimoto et al. (2022) and Eq. (6) gives its analytical explanation.
@@
Before calculating the variance, we first specify an initial state for the incomplete alternating state by considering how the alternating state is experimentally prepared. The experiment prepares the initial state by ramping up a strongly tilted double-well potential in one direction (see Sec. I. B. of the supplementary material of Ref. Wienand et al. (2023)). Thus, we expect that the experimental initial state is well described by a product state, assuming that the initial density matrix is approximated by
@@
We analytically derive the exact and asymptotic forms of the variance for the dynamics starting from the incomplete alternating state of Eq. (7). As detailed in Sec. V of the Supplemental Material SM , the two-point correlator becomes
@@
in the thermodynamic limit ($L\rightarrow\infty$). Here the density matrix is given by $\hat{\rho}(t)=e^{-{\rm i}\hat{H}t}\hat{\rho}_{\rm alt}e^{{\rm i}\hat{H}t}$ with Eq. (1) and $U=0$. Following the derivation given in Sec. VI of the Supplemental Material SM , we obtain
@@
with the function $\delta\sigma_{M}(t)^{2}\coloneqq M[n_{\rm even}(1-n_{\rm even})+n_{\rm odd}(1-n_{\rm odd})]/2+J_{0}(4t)[n_{\rm even}(1-n_{\rm even})-n_{\rm odd}(1-n_{\rm odd})][\sum_{m=0}^{M-1}(-1)^{m}]/2$. The time-dependent term of $\delta\sigma_{M}(t)^{2}$ is much smaller than the time-independent term when $M$ is large. Hence, we have $\delta\sigma_{M}(t)^{2}\simeq\sigma_{M}(0)^{2}$ for $M\gg 1$, and thus $\sigma_{\rm sub}(t)^{2}$ can be interpreted as the variance from which one subtracts its initial value. Applying the asymptotic analysis used in Eq. (6) to Eq. (9), we derive
@@
This result elucidates that the incompleteness of the alternating state substantially affects the coefficient of the variance, but the exponents of time are robust against it.
In addition to Eq. (10), we shall compare the numerical calculations in a finite system with the experiment. Our numerical calculation is implemented for $402L=$ 40, $\nu=0.44$, and $I=0.91$ with the initial state of Eq. (7). These concrete values are taken from Ref. Wienand et al. (2023) (see Sec. II. C. of the supplementary material of Ref. Wienand et al. (2023)). We adopt the Hamiltonian defined by
@@
with the open boundary condition ($\hat{a}_{L}=0$). Following the numerical simulation proposed in Ref. Wienand et al. (2023), we here add a random potential $V_{j}$, which is sampled from a uniform distribution with the range $[-\Delta,\Delta]$. Here, $\Delta\geq 0$ denotes the strength of the randomness. Under this setup, we investigate the variance $\sigma^{\prime}_{M}(t)^{2}\coloneqq\overline{\braket{\hat{N}_{M}^{\prime 2}}}-\left(~{}\overline{\braket{\hat{N}_{M}^{\prime}}}~{}\right)^{2}$ with $\hat{N}_{M}^{\prime}\coloneqq\sum_{m=-M/2}^{M/2-1}\hat{a}^{\dagger}_{m}\hat{a}_{m}$. Here, the overline $\overline{\bullet}$ denotes the ensemble average over the random potentials, and we use 10000 samples for this.
Figure 3: Comparison of our theoretical prediction with the experimental data. The parameters are $402L=$ 40, $\nu=0.44$, $I=0.91$, and $M=8$, which are taken from Ref. Wienand et al. (2023). The time evolution of the variance, denoted by circle and pentagon markers, is numerically obtained using Eq. (11) with $\Delta=0$ and $1$. For the comparison with the experiment, we plot the subtracted variance $2\sigma^{\prime}_{M}(t)^{2}-2\sigma^{\prime}_{M}(0)^{2}$ for the theoretical results and $\sigma_{\rm exp}(t)^{2}-\sigma_{\rm exp}(t_{0})^{2}$ for the experimental result because the experiment studies two-ladder lattice systems and thus the observed variance is twice as large as ours. Here, the experimental data $\sigma_{\rm exp}(t)^{2}$ is taken from Fig. S8 of Ref. Wienand et al. (2023), and $t_{0}$ is about $0.0006$. The dotted and dashed lines correspond to our analytical expressions of Eqs. (6) and (10) for $t\gg 1$, respectively. The plus marker denotes the numerical data of the 6th-order perturbative calculation.
Figure 3 compares our theoretical results with the experimental data of Ref. Wienand et al. (2023). Our numerical result with $\Delta=0$ can capture the variance growth in the early stage of the dynamics, but the deviation from the experimental data grows in time. On the other hand, when we turn on the random potential with $\Delta=1$, the numerical result well reproduces the experimental data. The saturation of the variance for $t\gtrsim 2$ is caused by the finite size effect and the disorder potential. These findings were reported in the numerical simulations of Ref. Wienand et al. (2023). We also implement the numerical perturbative calculation with $\Delta$, finding that the 6th-order perturbative result describes the time evolution before the saturation (see Sec. VII of the Supplemental Material for the detailed numerical method). These results show that the disorder effect is irrelevant in $t\lesssim 1$. For an arbitrary strength $\Delta$, we can identify such a time scale at which the disorder begins to affect the dynamics, by computing the variances for $\Delta=0$ and $\Delta\neq 0$.
The dotted and dashed lines in Fig. 3 show the analytical results for the linear growth of Eqs. (6) and (10). We find that the latter including the incompleteness of the initial alternating state exhibits the reasonable agreement with the experimental linear growth, while the former not including the incompleteness systematically deviates from the experimental data. We here emphasize that our analytical result (10) quantitatively describes the experimental data without any fitting parameters under the assumption of (7). Note that one can see that the disorder effect emerges before the finite $M$ effect by comparing the numerical data for $\Delta=0$ and 1 with Eq. (10) obtained in the limit ($L,M\rightarrow\infty$), though the localization length is larger than $M=8$ Wienand et al. (2023). In general, one can identify which the finite $M$ effect or the disorder effect emerges earlier, in the same manner with the help of Eq. (10).

### v3

@@ Abstract
Emergence of hydrodynamics in quantum many-body systems has recently garnered growing interest. The recent experiment of ultracold atoms [J. F. Wienand et al., Nat. Phys. (2024), doi:10.1038/s41567-024-02611-z] studied emergent hydrodynamics in hard-core bosons using a bipartite fluctuation, which quantifies how the particle number fluctuates in a subsystem. In this Letter, we theoretically study the variance of a bipartite fluctuation in one-dimensional noninteracting fermionic dynamics starting from an alternating state, deriving the exact solution of the variance and its asymptotic linear growth law for the long-time dynamics. To compare the theoretical prediction with the experiment, we generalize our exact solution by incorporating the incompleteness of the initial alternating state, deriving the general linear growth law analytically. We find that it shows good agreement with the experimentally observed variance growth without any fitting parameters. Furthermore, we estimate a time scale for the local equilibration using our exact solution, finding that the time scale is independent of the initial incompleteness. To investigate the interaction effect, we implement numerical studies for the variance growth in interacting fermions, which has yet to be explored experimentally. As a result, we find that the presence of interactions breaks the linear variance growth derived in the noninteracting fermions. Our exact solutions and numerical findings here lay a foundation for growing bipartite fluctuations in quantum many-body dynamics.
@@
Introduction.— Relaxation to an equilibrium state is ubiquitous in quantum many-body dynamics, bringing up fundamental and intriguing questions, e.g., how an isolated quantum system relaxes to a thermal equilibrium state Polkovnikov et al. (2011); Eisert et al. (2015); Nandkishore and Huse (2015); Torres-Herrera et al. (2015); Luca D’Alessio and Rigol (2016); Mori et al. (2018); Abanin et al. (2019); Gong and Hamazaki (2022). Over decades, such quantum thermalization has been intensively studied from broad perspectives, such as the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994), integrability Kinoshita et al. (2006); Gring et al. (2012); Cazalilla (2006); Rigol et al. (2006, 2007); Calabrese et al. (2011); Fagotti and Essler (2013); Langen et al. (2015); Ilievski et al. (2015); Tang et al. (2018), many-body localization Basko et al. (2006); Žnidarič et al. (2008); Pal and Huse (2010); Bardarson et al. (2012); Serbyn et al. (2013); Huse et al. (2014); Schreiber et al. (2015); Potter et al. (2015); Vasseur and Moore (2016); yoon Choi et al. (2016); Lüschen et al. (2017); Kohlert et al. (2019), and quantum scars Bernien et al. (2017); Shiraishi and Mori (2017); Turner et al. (2018a, b); Lin and Motrunich (2019); Shibata et al. (2020). To date, many experiments involving ultracold atoms and trapped ions have observed various phenomena related to quantum thermalization Kinoshita et al. (2006); Gring et al. (2012); Trotzky et al. (2012); Langen et al. (2013); Schreiber et al. (2015); Langen et al. (2015); Clos et al. (2016); Kaufman et al. (2016); Neill et al. (2016); Lüschen et al. (2017); Bernien et al. (2017); Tang et al. (2018); Kohlert et al. (2019).
@@
Figure 1: Schematic illustration of the main result. (a) One-dimensional fermions addressed in this work. The blue circles represent fermions, and they occupy every other site at the initial time. This state is referred to as an alternating state. The fermions hop to neighboring sites in time. The quantity of our interest is the variance of the particle number in a subsystem $A$ for a time-evolved state. (b) Schematic for the variance growth. The abscissa and ordinate, respectively, denote time $t$ and a subtracted variance, which is defined by the variance from which one subtracts its initial value. The two solid lines represent our analytical results, namely $2t/\pi$ and $2(n_{\rm even}-n_{\rm odd})^{2}t/\pi$, obtained by considering the variances growing from the complete and incomplete alternating states. Here, $n_{\rm even}$ and $n_{\rm odd}$ are averaged initial particle numbers at even and odd sites respectively, quantifying the incompleteness of the initial alternating state.
Recently, beyond the conventional quantum thermalization, hydrodynamic description based on local equilibrium has rapidly developed in quantum many-body systems. This situation is exemplified by the recent observation of electron fluids in strongly interacting systems Moll et al. (2016); Bandurin et al. (2016); Gooth et al. (2018); Krishna Kumar et al. (2017); Sulpizio et al. (2019); Aharon-Steinberg et al. (2022) and the success of generalized hydrodynamics being valid for integrable quantum models Castro-Alvaredo et al. (2016); Bertini et al. (2016); Bulchandani et al. (2017, 2018); Doyon and Yoshimura (2017); Doyon et al. (2017, 2018); Collura et al. (2018); De Nardis et al. (2018); Schemmer et al. (2019); Gopalakrishnan and Vasseur (2019); Doyon (2020); Alba et al. (2021); Malvania et al. (2021); Bouchoule and Dubail (2022); Essler (2023). When considering such hydrodynamic description, it is of essence to scrutinize how a system approaches a local equilibrium state. One of the most useful quantities for diagnosing local equilibrium is a bipartite fluctuation that quantifies how a physical quantity in a subsystem fluctuates. In a recent experiment of ultracold atoms for hard-core bosons Wienand et al. (2024), J. F. Wienand et al. utilized the bipartite fluctuation of the particle number in quench dynamics starting from the alternating state, studying the local equilibrium and the emergence of fluctuating hydrodynamics in the quantum many-body system. The experiment focused mainly on the hard-core bosons on a ladder system while also studying the ones on a one-dimensional system, which can be mapped into noninteracting fermions.
Despite the recent strong interest in emergent hydrodynamics in a quantum regime, understanding the experimentally observed bipartite fluctuation of Ref. Wienand et al. (2024) from the perspective of exact solutions is elusive even in the simplest one-dimensional hard-core bosons Groha et al. (2018); com , although several numerical works Fujimoto et al. (2020); Jin et al. (2020); Fujimoto et al. (2021, 2022); Cecile et al. (2024); Aditya and Roy (2024); Bhakuni and Lev (2023) exist. Hence, the fundamental understanding of such bipartite fluctuations is highly desirable now, and it is of immense importance to explain the recently observed bipartite fluctuation quantitatively via exact solutions.
@@
In this Letter, we consider one-dimensional fermions starting from the alternating state, theoretically studying time evolution for the variance of the particle numbers in a subsystem as schematically shown in Fig. 1 (a) and compare our theoretical prediction with the recent experimental data of Ref. Wienand et al. (2024). First, in the noninteracting fermions, we obtain the exact and simple expression of the variance and derive the asymptotic linear growth of the variance. Second, we compare our analytical result with the experiment of Ref. Wienand et al. (2024). For this purpose, we incorporate the incompleteness of the initial alternating state realized in the experiment into our exact solution, deriving the general linear growth law of the variance. This general law agrees well with the variance growth observed in Ref. Wienand et al. (2024) without any fitting parameters. The impact of the incompleteness is schematically displayed in Fig. 1 (b). Furthermore, utilizing the exact solution before taking the large-size limit, we analytically estimate a time scale for the local equilibration. Finally, we numerically investigate interaction effects on the variance growth using the time-evolving block decimation (TEBD) method Vidal (2003, 2004); Schollwöck (2011); Paeckel et al. (2019).
@@
Before ending the introduction, it is worth noting that a bipartite fluctuation has been studied in terms of an integrated current and full counting statistics. In the former case, most of previous works consider dynamics starting from a domain-wall initial state, investigating a bipartite fluctuation for a conserved quantity by connecting it to an integrated current via a continuity equation Antal et al. (2008); Eisler and Rácz (2013); Ljubotina et al. ; Moriya et al. (2019); Gamayun et al. (2020); Jin et al. (2021); Gopalakrishnan and Vasseur (2023); Wei et al. (2022); Krajnik et al. (2022a, b, 2024a, 2024b). One of the notable results is the logarithmic growth of the variance in noninteracting fermions Antal et al. (2008). In the latter case, a bipartite fluctuation has been employed to study static and dynamic properties of entanglement entropy Klich and Levitov (2009); Song et al. (2010, 2012); Rachel et al. (2012); Parez et al. (2021); Oshima and Fuji (2023); Bertini et al. (2023a) and fluctuations of physical quantities Schönhammer (2007); Collura et al. (2017); Stéphan and Pollmann (2017); Najafi and Rajabpour (2017); Arzamasovs and Gangardt (2019); Calabrese et al. (2020); Ares et al. (2021); Smith et al. (2021); McCulloch et al. (2023); Hercé et al. (2023); Bertini et al. (2023b).
@@
with the particle-number operator $\hat{n}_{j}\coloneqq\hat{a}_{j}^{\dagger}\hat{a}_{j}$ and the interaction parameter $U$. Here, the boundary condition is basically a periodic boundary condition ($\hat{a}_{L}=\hat{a}_{-L}$), but we use an open boundary condition ($\hat{a}_{L}=0$) in several cases. In the latter case, we will explicitly specify the boundary condition. The initial state is the alternating state defined by |$\ket{\psi_{\rm alt}}\coloneqq\prod_{j=-L/2}^{L/2-1}\hat{a}_{2j+1}^{\dagger}\ket{0}$. We denote a quantum state at time $t$ by |$\ket{\psi(t)}$ and assume that the state obeys the Schrödinger equation ${\rm i}d\ket{\psi(t)}/dt=\hat{H}\ket{\psi(t)}$. Here, the Dirac constant $\hbar$ is set to be unity. Note that our model with $U=0$ is equivalent to 1D hard-core bosons. Thus, our theoretical results based on this model are relevant to experimental results for 1D hard-core bosons of Ref. Wienand et al. (2024).
In this Letter, we shall study the fluctuation of the particle number in a subsystem. The operator for the bipartite particle-number is defined by $\hat{N}_{M}\coloneqq\sum_{j=0}^{M-1}\hat{a}_{j}^{\dagger}\hat{a}_{j}$ with a positive integer $M$. Then, a generating function for the $n$th moment of $\hat{N}_{M}$ is given by $G_{M}(\lambda,t)\coloneqq\braket{e^{\lambda\hat{N}_{M}}}_{t}$, where we introduce the notation $\braket{\bullet}_{t}\coloneqq{\rm Tr}[\hat{\rho}(t)\bullet]$ with the density matrix $\hat{\rho}(t)\coloneqq\ket{\psi(t)}\bra{\psi(t)}$ |. We can compute the moments by differentiating $G_{M}(\lambda,t)$ with respect to $\lambda$.
Exact solution for the variance of the noninteracting fermions.— We study the time evolution for the variance of $\hat{N}_{M}$ in the noninteracting fermions ($U=0$) under the thermodynamic limit ($L\rightarrow\infty$). As described in Sec. I of the Supplemental Material (SM) SM , the generating function $G_{M}(\lambda,t)$ becomes Eisler and Rácz (2013); Schönhammer (2007); Parez et al. (2021)
@@
The detailed derivation of Eq. (4) is given in Sec. II of SM SM . The next task is to take the limit $M\rightarrow\infty$ in Eq. (4). As proved in Sec. III of SM SM , we have $\lim_{M\rightarrow\infty}M\left(1-J_{0}(4t)^{2}-2\sum_{k=1}^{M-1}J_{k}(4t)^{2}\right)/4=0$ for $t>0$ and $\lim_{M\rightarrow\infty}\sum_{k=1}^{M-1}kJ_{k}(4t)^{2}/2=4t^{2}\left(J_{0}(4t)^{2}+J_{1}(4t)^{2}\right)-tJ_{0}(4t)J_{1}(4t)$. We utilize these formulae and Eq. (4), deriving the following exact and simple expression of the variance $\sigma(t)^{2}\coloneqq\lim_{M\rightarrow\infty}\sigma_{M}(t)^{2}$,
@@
The detailed derivation of Eq. (6) is given in Sec. IV of SM SM . The essential consequence of Eq. (6) is that the variance is proportional to time $t$ for $t\gg 1$. This linear growth law is entirely different from the variance growth in noninteracting fermionic dynamics starting from the domain-wall state |$\ket{\psi_{\rm DW}}\coloneqq\prod_{j\in\mathbb{Z}_{<0}}\hat{a}_{j}^{\dagger}\ket{0}$, where the variance was shown to obey the logarithmic growth Antal et al. (2008); Moriya et al. (2019). Note that the linear growth law was numerically confirmed in Ref. Fujimoto et al. (2022) and Eq. (6) gives its analytical explanation.
@@
Before calculating the variance, we first specify an initial state for the incomplete alternating state by considering how the alternating state is experimentally prepared. The experiment prepares the initial state by ramping up a strongly tilted double-well potential in one direction (see Sec. I. B. of the supplementary information of Ref. Wienand et al. (2024)). Thus, we expect that the experimental initial state is well described by a product state, assuming that the initial density matrix is approximated by
@@
We analytically derive the exact and asymptotic forms of the variance for the dynamics starting from the incomplete alternating state of Eq. (7). As detailed in Sec. V of SM SM , the two-point correlator becomes
@@
in the thermodynamic limit ($L\rightarrow\infty$). Here the density matrix is given by $\hat{\rho}(t)=e^{-{\rm i}\hat{H}t}\hat{\rho}_{\rm alt}e^{{\rm i}\hat{H}t}$ with Eq. (1) and $U=0$. Following the derivation given in Sec. VI of SM SM , we obtain the variance as
$\sigma_{M}(t)^{2}=\delta\sigma_{M}(t)^{2}+(n_{\rm even}-n_{\rm odd})^{2}\left[\dfrac{M}{4}\left(1-J_{0}(4t)^{2}-2\sum_{m=1}^{M-1}J_{m}(4t)^{2}\right)+\dfrac{1}{2}\sum_{m=1}^{M-1}mJ_{m}(4t)^{2}\right].$ (9)
with the function $\delta\sigma_{M}(t)^{2}\coloneqq M[n_{\rm even}(1-n_{\rm even})+n_{\rm odd}(1-n_{\rm odd})]/2+J_{0}(4t)[n_{\rm even}(1-n_{\rm even})-n_{\rm odd}(1-n_{\rm odd})][\sum_{m=0}^{M-1}(-1)^{m}]/2$. Under the setup, we can derive
@@
Here, the time-dependent term of $\delta\sigma_{M}(t)^{2}$ is much smaller than the time-independent term when $M$ is large. Hence, we have $\delta\sigma_{M}(t)^{2}\simeq\sigma_{M}(0)^{2}$ for $M\gg 1$. Thus $\sigma_{\rm sub}(t)^{2}$ can be interpreted as the variance from which one subtracts its initial value. Applying the asymptotic analysis used in Eq. (6) to Eq. (10), we derive
@@
This result elucidates that the incompleteness of the alternating state substantially affects the coefficient of the variance, but the exponents of time are robust.
Figure 3: Comparison of our theoretical prediction with the experimental data. The parameters are $402L=$ 40, $\nu=0.44$, $I=0.91$, and $M=8$, which are taken from Ref. Wienand et al. (2024). The time evolution of the variance, denoted by circle and pentagon markers, is numerically obtained using Eq. (12) with $\Delta=0$ and $1$. For the comparison with the experiment, we plot the subtracted variance $2\sigma^{\prime}_{M}(t)^{2}-2\sigma^{\prime}_{M}(0)^{2}$ for the theoretical results and $\sigma_{\rm exp}(t)^{2}-\sigma_{\rm exp}(t_{0})^{2}$ for the experimental result because the experiment studies two-ladder lattice systems and thus the observed variance is twice as large as ours. Here, the experimental data $\sigma_{\rm exp}(t)^{2}$ is taken from Fig. S8 of Ref. Wienand et al. (2024), and $t_{0}$ is $0.0006031857894892$. The dotted and dashed lines correspond to our analytical expressions of Eqs. (6) and (11) for $t\gg 1$, respectively. The plus marker denotes the numerical data of the 6th-order perturbative calculation.
In addition to Eq. (11), we shall compare the numerical calculations in a finite system with the experiment. Our numerical calculation is implemented for $402L=$ 40, $\nu=0.44$, and $I=0.91$, which are taken from Ref. Wienand et al. (2024) (see Sec. II. C. of the supplementary information of Ref. Wienand et al. (2024)). We adopt the Hamiltonian defined by
@@
with the open boundary condition ($\hat{a}_{L}=0$). Following the numerical simulation in Ref. Wienand et al. (2024), we here add a random potential $V_{j}$ sampled from a uniform distribution with the range $[-\Delta,\Delta]$. Here, $\Delta\geq 0$ denotes the strength of the randomness. Under this setup, we investigate the variance $\sigma^{\prime}_{M}(t)^{2}\coloneqq\overline{\braket{\hat{N}_{M}^{\prime 2}}}-\left(~{}\overline{\braket{\hat{N}_{M}^{\prime}}}~{}\right)^{2}$ with $\hat{N}_{M}^{\prime}\coloneqq\sum_{m=-M/2}^{M/2-1}\hat{a}^{\dagger}_{m}\hat{a}_{m}$. Here, the overline $\overline{\bullet}$ denotes the ensemble average over the random potentials, and we use 10000 samples for this.